OpenStudy (anonymous):

Please Help me!!!!!! ~ Select the equation of the line that passes through the point (2, –1) and has slope 3 in point-slope form

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(2\quad ,&-1) \end{array} \\\quad \\ slope = m= 3 \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"} \)

OpenStudy (anonymous):

Okay i puged it n now what im still confused.. (O.o)